# Can Large Language Models Act as Ensembler for Multi-GNNs?

### v1

@@ 5. EXPERIMENT
Model type Model Cora (Acc/%) PubMed (Acc/%) ogbn-arXiv (Acc/%)
MLP MLP 66.42 82.40 61.47
GNN GCN 85.97 83.78 70.81
GAT 86.71 83.59 70.05
GIN 85.60 82.03 67.03
GraphSAGE 84.87 87.01 70.35
Graphormer 80.41 88.75 72.81
GT 86.42 88.24 -
CoarFormer 88.69 89.75 71.66
Language Model BERT 80.15 93.91 72.78
SentenceBERT 78.82 92.49 71.42
DeBERTa 77.79 93.45 72.90
LLM Based DGTL 81.10 87.10 -
@@
LensGNN-[GCN+GAT] 88.56 94.11 75.78
LensGNN-[GCN+GIN] 91.88 94.47 75.89
LensGNN-[GAT+GIN] 88.19 95.43 74.31
LensGNN-ALL 90.40 95.68 75.91
5.1. Experimental Settings
Datasets. We use three benchmark datasets whose details are given as follows. Statistics of these datasets are summarized in Table 1.
Cora is a citation network containing 2,708 research papers. These papers are categorized into seven classes. In this graph, each node represents a scientific paper, and each edge represents a citation. Bag-of-words representation is used as the feature vector for each node, which is also assigned a label indicating its research field. The task for this dataset is node classification.
PubMed is a citation network dataset containing 19,717 research papers from the biomedical field. These papers are categorized into three classes, and each paper is represented by a bag-of-words feature vector. In this graph, each node represents a scientific paper, and each edge represents a citation. The task associated with this dataset is node classification.
ogbn-arXiv is part of the Open Graph Benchmark (OGB) and contains 169,343 academic papers scraped from arXiv. In this dataset, the nodes represent arXiv papers, and the edges represent citation relationships. These papers are categorized into 40 subject areas. We construct node attributes using the title and abstract of the papers. The task associated with this dataset is node classification.
Baselines. We compare LensGNN with 17 baselines, which can be categorized into four types. The first type involves (1) MLP. The second type involves (2) GCN (Kipf and Welling, 2016), (3) GAT (Veličković et al., 2017), (4) GIN (Xu et al., 2019), (5) GraphSAGE (Hamilton et al., 2017a), (6) Graphormer (Ying et al., 2021), (7) GT (Dwivedi and Bresson, 2012) and (8) CoarFormer (Kuang et al., 2024) seven popular GNNs for graph tasks. For fairness, we use the node representations obtained from a pre-trained LM as input to the GNN models for learning node classification. The third type involves (9) BERT (Devlin, 2018), (10) SentenceBERT (Reimers, 2019) and (11) DeBERTa (He et al., 2020) three language models. They are directly trained on graph datasets with the raw text. The fourth type involves (12) DGTL (Qin et al., 2024), (13) SNS-GPT4 (Li et al., 2024), (14) InstructGraph-INS (Wang et al., 2024), (15) InstructGLM (Ye et al., 2024), (16) GAugLLM (Fang et al., 2024) and (17) Baichuan2 (Yang et al., 2023) six LLM-based methods designed for graph tasks. We report the results from the original papers due to limitations in computational power. For Baichuan2, we directly used it for node class prediction without fine-tuning. Due to the space limitation, we move details on baselines to Appendix A.
Setup. We implement and conduct experiments for LensGNN by PyTorch. We primarily use Baichuan2-13B (Yang et al., 2023) as the backbone LLM for LensGNN. LensGNN ensemble multiple GNN models, including GCN (Kipf and Welling, 2016), GAT (Veličković et al., 2017), GIN (Xu et al., 2019), each with two layers. We perform a grid search to fine-tune hyperparameters based on the results on the validation set. The specific hyperparameter search space is shown in the Table 2. Some fixed LoRA settings included capping each training sample at 2,047 tokens and using half-precision (FP16) for LoRA fine-tuning, with a batch size of 4 per GPU and gradient updates every step. We utilize a cosine-type learning rate scheduler and set the warmup ratio to 0.1. Specifically, we utilize LoRA+ (Hayou et al., 2024). For baselines that reported results on the datasets, we directly report the results from their original papers in the table. Due to computational power limitations, for those datasets where the original papers of the baselines did not report results, we left them blank in the table. We run all the experiments on a server with a single NVIDIA Tesla A100 GPU.
Table 4. Ablation study including ‘GNN Encoder’, ‘Alignment’, ‘With Text’ and ‘With Neighbor’.
[…]
These elements enhance the model’s ability to integrate representations and contextual information, resulting in higher accuracy across various datasets. GNN Encoder Alignment With Text With Neighbor Cora (Acc/%) PubMed (Acc/%) ogbn-arXiv (Acc/%) - - Yes 1 87.82 93.77 74.47 GCN - Yes 1 89.29 94.23 74.35 GAT - Yes 1 90.03 94.26 73.37 GIN - Yes 1 88.92 94.82 73.26 GCN, GAT, GIN Yes No 1 19.92 40.06 34.92 GCN, GAT, GIN No Yes 1 90.77 93.96 73.73 GCN, GAT, GIN Yes Yes 0 84.87 94.21 72.03 GCN, GAT, GIN Yes Yes 1 90.40 95.68 75.91
[…]
@@ 5.1. Experimental Settings
5.2. Overall Performance Comparison (RQ1)
We conduct experiments to compare LensGNN with 17 state-of-the-art competitors on Cora, PubMed and ogbn-arXiv datasets. For Cora and PubMed, we randomly split nodes into 60%, 20%, and 20% for training, validation and testing, and measure the performance of all models on the test set. For ogbn-arXiv, we split the dataset as suggested in (Hu et al., 2020). We use pre-trained LM to encode node representations based on their original text for all GNN models. For fairness, we use the same pre-trained LM, SentenceBERT (Reimers, 2019), as in LensGNN. Additionally, we compare our backbone LLMs, Baichuan2 (Yang et al., 2023), on the node classification task, where we use Baichuan2 for node label inference directly without fine-tuning. For LensGNN, we report the results of several variants of LensGNN , including a model that integrates GCN, GAT, and GIN simultaneously, as well as variants that ensemble each pairwise combination of these three GNNs.
Table 3 shows the classification accuracy of all the methods. From the tables, we make the following observations: (1) LensGNN outperforms the competitive methods over all the datasets. (2) Under the condition of using the same pre-trained LM to encode node representations, LensGNN outperforms individual GNN models’ classification performance across all datasets. Particularly on the PubMed dataset, LensGNN’s classification accuracy surpasses that of the best GNN methods by more than 5%. This demonstrates the effectiveness of using LLMs to ensemble multiple GNN models in node classification tasks. (3) The LM-based methods encode node representations using raw text to achieve classification. However, these approaches discard the structural properties unique to graph data, resulting in the loss of valuable information and an inability to achieve optimal performance. This indicates the necessity of structural information in graph-related tasks. (4) Some methods attempt to use LLMs to accomplish various graph tasks. For example, InstructGLM describes graph structures in natural language and feeds them into a LLM to achieve node classification. However, LensGNN still achieved better performance across all datasets because LensGNN can leverage the LLM to ensemble the strength of multiple GNNs. (5) Baichuan2 showed poor performance and could not even outperform the language models. This is because LLMs that have not been fine-tuned are not suitable for node classification tasks on graph data. Overall, LensGNN harnesses the strengths of multiple GNNs through LLMs and avoids the tedious effort of cherry-picking GNNs across datasets, thereby achieving the best performance across all datasets.
5.3. Ablation Study (RQ2)
Figure 4. Inference Speed and Accuracy of Our Method Based on Different Backbone LLMs.
In this section, we conduct an ablation study on the Cora, PubMed, and ogbn-arXiv datasets to evaluate the impact of each model component. Specifically, we explore various configurations of the ‘GNN Encoder’ ( using different GNNs ), ‘Alignment’ ( whether multiple GNNs are aligned when training GNNs ), ‘With Text’ ( whether node text is included in prompt) and ‘With Neighbor’ ( the number of neighbour hops of the node text when inputting LLMs ). The results are presented in Table 4. From the table, we have the following observations:
Performance of different GNN encoders. We find that ensembling multiple GNNs noticeably outperforms using a single GNN. For example, on the ogbn-arxiv dataset, the best result from a single GNN is 74.35%, while LensGNN achieves 75.91%. This clearly illustrates that by integrating the strengths of various types of GNNs, LensGNN can effectively enhance overall performance.
Impact of Alignment. On the PubMed dataset, the results for aligned GNN and unaligned GNN are 93.96% and 95.68%, respectively, while on the ogbn-arXiv dataset, the results are 73.73% and 75.91%, respectively. This indicates that aligning multiple GNNs representation space can help LLMs better understand graph tokens. Although the performance of aligned GNN is slightly lower than unaligned GNN on the Cora dataset, this may be due to the insufficient number of training samples.
The Importance of Node Text. It can be observed that node text plays a key role in the supervised fine-tuning of LLMs. Through these node texts, the model can achieve a deeper understanding of semantics, resulting in outstanding accuracy.
Role of Neighbors. The textual information of neighbors provides support for LLM’s understanding of the semantic information of nodes, enabling LLM grasp structural information more deeply, thereby improving classification performance.
Overall, the ablation study confirms that each component of the model contributes significantly to its overall performance.
5.4. Performance with different LLMs (RQ3)
To further demonstrate the effectiveness of LensGNN, we use other LLMs as backbones, including three smaller LMs: BERT-base (Devlin, 2018), T5-base (Raffel et al., 2023), the encoder-only variant of T5-base, and two other LLMs: Falcon-7B (Almazrouei et al., 2023) and InternLM2.5-7B-chat (Cai et al., 2024). The experiments based on these LLMs include aligning each language model separately with the representation of each type of GNN, as well as aligning them simultaneously with all three types of GNNs (GCN, GAT, GIN) and performing an ensemble. All results are presented in Table 5.
From the table, we make the following observations: (1) Smaller LMs BERT-base, T5-base, and T5-base (Encoder only) can achieve good results when ensemble a single GNN, but the performance drops significantly when ensemble multiple GNN models. This suggests that models with a smaller parameters lack the capability to integrate multiple GNNs. (2) In contrast, LLMs with large parameters that ensemble a single GNN generally perform better than smaller LMs in most cases. (3) Across different large parameter language models, the ensemble of multiple GNN models achieved the best performance on all datasets. Overall, all the observations indicate that LLMs can better achieve the ensemble of GNN models, and the multi-GNNs ensemble can harnesses the strengths of multiple GNNs to achieve the best predictions.
5.5. Model Efficiency Study (RQ4)
In this section, we analyze the computational efficiency of LensGNN during both the training and inference phases.
@@ 5.5. Model Efficiency Study (RQ4)
Model Training Efficiency. We divide the process of LensGNN into two phases: Aligning multi-GNNs and Ensembling multi-GNNs with LLM. In the first phase, the training efficiency primarily depends on the efficiency of the GNN encoders utilized, while also being related to the hyperparameter Graph token $t$. In the second phase, the training efficiency predominantly relies on the fine-tuning efficiency of the LLMs. We employ LoRA technology to reduce the number of parameters requiring fine-tuning. For instance, in the case of Baichuan2-13B, the trainable parameters constitute only 0.0470% of the total parameters. This allows us to fine-tune the 13B parameter LLM in a single 80G Nvidia A100 environment, enabling a training batch size of 4 per device.
@@
Model Inference Efficiency. During inference, all parameters are frozen. Since the model parameters in the first stage are much smaller than the LLM in the second stage, the time cost in the inference phase mainly comes from the second stage’s LLM inference. In our experiments, with a maximum input length of 2047 tokens, the inference speed and accuracy of our method based on Baichuan2-13B, Falcon-7B, and InternLM2.5-7B-chat on the Cora and Pubmed datasets are illustrated in Figure 4. Overall, the inference speed ranges from 2 to 6 samples per second. Although a higher parameter count in the LLM results in reduced inference speed, it concurrently yields superior accuracy. In practical deployments, the choice of backbone LLM can be made by considering the balance between computational power and application requirements.
5.6. Hyperparameter Sensitivity Analysis (RQ5)
The number of Graph Tokens $t$ is an important hyperparameter, representing how many tokens each node is divided into before being input into LLM. We perform experiments on the Cora and Pubmed dataset with $t$ set to values of 1, 2, 4, 8, 16 and 32, and the results are shown in Figure 5.
@@ 5.6. Hyperparameter Sensitivity Analysis (RQ5)
The results indicate that as the number of Graph Tokens increases, the model’s performance initially rises before gradually declining. The optimal experimental results are achieved when the number of Graph Tokens is set to 8 on both datasets. When the number of tokens is too small, the distinction between nodes becomes insufficient, making it difficult for LLMs to fully learn the semantic information. On the other hand, when there are too many tokens, the representation of Graph Tokens may introduce noise or redundant information, which interferes with the model’s understanding of the graph representations. Therefore, an appropriate number of tokens helps LLMs effectively capture more complicated structural information, enhancing semantic comprehension and leading to better performance.
[…]
6. CONCLUSION
@@
Appendix A Baselines

### v2

@@ 5 EXPERIMENT
Model type Model Cora PubMed ogbn-arXiv Citeseer Wiki-CS
MLP MLP 66.42 82.40 61.47 71.13 68.41
GNN GCN 85.97 83.78 70.81 77.74 77.15
GAT 86.71 83.59 70.05 78.21 78.94
GIN 85.60 82.03 67.03 75.07 68.97
GraphSAGE 84.87 87.01 70.35 73.61 79.56
Graphormer 80.41 88.75 72.81 71.28 72.07
Language Model BERT 80.15 93.91 72.78 73.17 78.33
SentenceBERT 78.82 92.49 71.42 72.79 77.92
DeBERTa 77.79 93.45 72.90 73.13 75.11
LLM Ensembler LensGNN-[GCN+GAT] 88.56 94.11 75.78 78.05 18.41
LensGNN-[GCN+GIN] 91.88 94.47 75.89 76.64 83.47
LensGNN-[GAT+GIN] 88.19 95.43 74.31 78.99 82.98
LensGNN-ALL 90.40 95.68 75.91 79.31 81.78
5.1 Experimental settings
Datasets. We use eight benchmark datasets, comprising five node classification datasets: Cora, PubMed, ogbn-arXiv, Citeseer and Wiki-CS, and three molecular graph classification datasets Wu et al. (2018): BACE, BBBP and ClinTox. Details on these datasets are provided in the Appendix A.
Baselines. We compare our method with 19 baselines: MLP, GNN models, LM-based models and LLM-based models. Specifically, GNN models include GCN Kipf and Welling (2016), GAT Veličković et al. (2017), GIN Xu et al. (2019), GraphSAGE Hamilton et al. (2017a) and Graphormer Ying et al. (2021). LM-based models consist of BERT Devlin (2018), SentenceBERT Reimers (2019) and DeBERTa He et al. (2020). LLM-based models include DGTL Qin et al. (2024), SNS-GPT4 Li et al. (2024), GAugLLM Fang et al. (2024) and Baichuan2-13B Yang et al. (2023), which all employ 13B LLMs. We also compare LensGNN with other 7B-LLM-based models: GraphGPT Tang et al. (2024), LLaGA Chen et al. (2024), OFA Liu et al. (2024) and GOFA Kong et al. (2024). Due to the space limitation, details on these baselines are given in Appendix B.
Setup. We use Baichuan2 Yang et al. (2023) and InternLM2.5 Cai et al. (2024) as the backbone LLM for LensGNN. In the experiments, we ensemble three widely adopted GNN models: GCN, GAT and GIN, each with two layers. For GNN models, we use node representations obtained from the pre-trained SentenceBERT Reimers (2019) as input. For more details on experiment setup, refer to Appendix C.
5.2 Node classification results
Since the large number of baselines report their results on different datasets and settings, we show the classification results in three tables, respectively. Table 1 compares LensGNN with MLP, GNN models and small language models. Tables 2 summarizes the results against LLM-based models. Table 3 evaluates the performance of LensGNN in the 20-shot setting. From these tables, we observe:
(1) LensGNN consistently outperforms all the GNN models and LM-based methods across all the datasets. For GNN models, while they initialize node features with SentenceBERT to capture text semantics, LLMs have much stronger capability in text understanding and generation, which explains the advantage of LensGNN. For LM-based models, in addition to the weak expressiveness of small language models, they ignore the rich structural information of graphs, degrading their performance.
(2) LensGNN achieves better performance than LLM-based models, even in the 20-shot setting. In particular, we see that Baichuan2-13B performs very poorly. We speculate this is because LLMs that have not been fine-tuned are not suitable for node classification tasks on graph data. Different from other LLM-based models, LensGNN integrates the strengths of multiple GNNs and employs two-phases of alignment, which explains its superior performance.
Table 2: Comparison on classification accuracy (%) with LLM-based models. * denotes the best variant of our model and - means that the results are missing from their original papers.
Model Cora PubMed ogbn-arXiv
GraphGPT-MIX-7B - 74.16 64.76
LLaGA-HO-7B(GENERAL) - 94.45 75.01
InstructGLM-Llama-7B 87.08 93.84 75.70
DGTL 81.10 87.10 -
@@ 5.2 Node classification results
LensGNN*-InternLM2.5-7B 90.03 95.13 74.24
LensGNN*-Baichuan2-7B 89.85 95.08 75.88
LensGNN*-Baichuan2-13B 91.88 95.68 75.91
Table 3: Classification accuracy in 20-shot setting.
Model Cora PubMed
GCN 73.5 68.0
GAT 72.8 68.1
GIN 75.7 69.3
OFA 75.34 77.89
GOFA 77.08 87.33
LensGNN*-Baichuan2-7B 78.09 89.19
LensGNN*-InternLM2.5-7B 79.93 87.75
5.3 Graph classification results
Table 4: Graph classification results.
Model BACE BBBP ClinTox
ACC AUC ACC AUC ACC AUC
GCN 74.01 84.92 84.14 79.00 90.93 87.23
GAT 74.67 80.85 76.34 66.12 91.61 86.37
GIN 68.09 85.66 83.65 79.02 86.91 85.80
LensGNN-[GCN+GAT] 73.02 72.67 86.09 81.29 98.99 94.00
LensGNN-[GCN+GIN] 75.32 74.59 88.04 83.27 99.32 96.00
LensGNN-[GAT+GIN] 76.31 75.99 86.09 81.64 98.99 95.81
LensGNN-ALL 80.59 80.33 88.53 85.37 98.99 97.63
We further evaluate the effectiveness of LensGNN on the graph classification task. We use classification accuracy (ACC) and AUC scores as the evaluation metrics. The results are given in Table  4. From the table, LensGNN outperforms GNN baselines in most cases, which shows that combining GNNs and LLMs is inspiring. Further, LensGNN-ALL generally performs better than its variants. This verifies that leveraging strengths of multiple GNNs is useful. Surprisingly, we find that on BACE, LensGNN and its variants achieve smaller AUC scores than GNN models. We step into the datasets and observe that BACE is a label-balanced dataset while BBBP and ClinTox are highly imbalanced. Note that, for balanced dataset, ACC is more convincing than AUC score, while AUC score is a better indicator than ACC on imbalanced dataset. Therefore, the ACC advantage of LensGNN on BACE and the AUC leads on other two datasets have evidently demonstrated the superiority of our method.
5.4 Ablation study
We next systematically conduct an ablation study on Cora, PubMed, and ogbn-arXiv datasets to evaluate the importance of major model components. Specifically, we explore various configurations of “GNN Encoder” (using different GNNs ), “Alignment” (whether multiple GNNs are aligned when training GNNs ), “With Text” (whether node text is included in prompt) and “With Neighbor” (whether neighborhood information is included or not). The results are presented in Table 5.
Performance of different GNN encoders. Ensembling multiple GNNs noticeably outperforms using a single GNN. For example, on the ogbn-arxiv dataset, the best result from a single GNN is 74.35%, while LensGNN achieves 75.91%. This illustrates that by integrating the strengths of various GNNs, LensGNN can effectively enhance the overall performance.
Impact of Alignment. On the PubMed dataset, the results for unaligned GNN and aligned GNN are 93.96% and 95.68%, respectively, while on the ogbn-arXiv dataset, the results are 73.73% and 75.91%, respectively. This shows that multi-GNN alignment can help LLMs better understand graph tokens. Although the performance of aligned GNN is slightly lower than unaligned GNN on the Cora dataset, this may be due to the insufficient number of training samples.
Importance of Node Text. It can be observed that node text plays a key role in the supervised fine-tuning of LLMs. Through these node texts, the model can achieve a deeper understanding of semantics, resulting in outstanding accuracy.
Role of Neighbors. The textual information of neighbors provides support for LLM’s understanding of the semantic information of nodes, thereby improving classification performance.
Table 5: Ablation study results.
GNN Encoder Alignment With Text With Neighbor Cora (Acc/%) PubMed (Acc/%) ogbn-arXiv (Acc/%)
- - Yes 1 87.82 93.77 74.47
GCN - Yes 1 89.29 94.23 74.35
GAT - Yes 1 90.03 94.26 73.37
GIN - Yes 1 88.92 94.82 73.26
GCN, GAT, GIN Yes No 1 19.92 40.06 34.92
GCN, GAT, GIN No Yes 1 90.77 93.96 73.73
GCN, GAT, GIN Yes Yes 0 84.87 94.21 72.03
GCN, GAT, GIN Yes Yes 1 90.40 95.68 75.91
5.5 Performance with different backbones
Table 6: The classification accuracy (%) with different LLMs on Cora and PubMed.
@@ 5.5 Performance with different backbones
To study the impact of backbone models on LensGNN, we use different backbones, including three small LMs: BERT-base Devlin (2018), T5-base Raffel et al. (2023), the encoder-only variant of T5-base, and two other LLMs: Falcon-7B Almazrouei et al. (2023) and InternLM2.5-7B-chat Cai et al. (2024). All the results are presented in Table 6. From the table, we see that
(1) Small LMs BERT-base, T5-base, and T5-base (Encoder only) perform well when applied to a single GNN, but the performance drops significantly when ensembling multiple GNN models. This suggests that models with less parameters lack the capability to integrate multiple GNNs.
(2) For LLMs, they generally perform better than LM backbones. For example, for GAT on Cora, the best result for LM backbones is 81.85%, while that for LLM backbones is 90.03%. This is because LLMs can provide richer text semantics. Further, compared with single GNN, the ensemble of multiple GNNs with LLMs leads to better performance, which verifies that unifying the strengths of multiple GNNs is beneficial for node classification.
5.6 Model efficiency study
@@
Training Efficiency. The training of LensGNN includes two main phases: aligning multi-GNNs and ensembling multi-GNNs with LLM. In the first phase, the time complexity depends on the GNN encoders. Since the used GNNs have a linear time complexity w.r.t. the number of nodes in the graph, the training process is efficient. In the second phase, fine-tuning LLMs is the main computational cost. We employ LoRA to reduce the number of parameters to be fine-tuned. For example, for Baichuan2-13B, the trainable parameters constitute only 0.0470% of the total parameters. This allows us to fine-tune the 13B parameter LLM in a single 80G Nvidia A100 environment. Compared to GraphGPT Tang et al. (2024), the increase of our training computational cost primarily stems from employing more than one GNN. However, in contrast to LLM, the number of parameters in GNN is considerably smaller, which increases marginal training cost.
@@
Inference Efficiency. During inference, all the parameters are frozen. The major cost of model inference comes from LLMs. In our experiments, with a maximum input length of 2047 tokens, the inference speed and accuracy of our method based on Baichuan2-13B, Falcon-7B, and InternLM2.5-7B-chat on the Pubmed dataset is illustrated in Figure 3. Overall, the inference speed ranges from $2$ to $4$ samples per second. Although a larger LLM results in reduced inference speed, it concurrently yields superior accuracy.
Figure 3: Inference speed and accuracy comparison.
5.7 Hyperparameter sensitivity analysis
@@
We end this section with model sensitivity analysis on the number of graph tokens $t$, which represents how many tokens each node is mapped into before fed into LLM. We conduct experiments on Cora and Pubmed with varying $t$ values, and the results are shown in Figure 4. From the figure, we see that, as the number of graph tokens increases, the model’s performance first rises and then decreases. For both datasets, the best results are achieved when $t=8$. When $t$ is small, graph token representations cannot well capture the semantic information. On the other hand, when $t$ is large, the representation of graph tokens could be noise-corrupted, which adversely affects the model’s understanding on graph structure. Therefore, an appropriate number of tokens helps LLMs effectively capture structural information, enhancing semantic comprehension and leading to better performance.
Figure 4: Hyperparameter sensitivity analysis.
6 CONCLUSION
@@
7 LIMITATIONS
Our method could suffer from several limitations:
First, due to limited computational resources, we do not conduct experiments on LLMs with parameters exceeding 13B. This implies that the upper limit of our method’s capabilities has not been fully investigated. For larger-scale LLMs that are utilized in production environments, they could further improve our model’s performance.
Second, our method relies on manually crafted prompts to ensemble multi-GNNs. Future studies on prompt design could be a promising direction.
Third, we only evaluated the performance based on node classification and graph classification tasks. Evaluation on other tasks, such as graph dataset generation and graph task interpretability analysis, are valuable research questions.
8 ETHICAL STATEMENT
Our work falls under basic research and is not tied to specific applications; therefore, whether our method will be abused and cause negative social impacts depends on the specific applications in which others use our method. In addition, our work does not involve any stakeholders benefiting or being disadvantaged, nor does it involve vulnerable groups. The datasets we used are all commonly used public datasets, posing no privacy risks, and aligned with their intention for scientific research. For the datasets, pre-trained models, and training frameworks utilized, we adhered to their respective open-source licenses: community license for Baichuan-2 Yang et al. (2023), Apache License 2.0 for InternLM2.5 Cai et al. (2024) , Falcon-7B Almazrouei et al. (2023) and LLaMA-Factory Zheng et al. (2024)
Appendix A Datasets
We use eight benchmark datasets, comprising five node classification datasets: Cora, PubMed, ogbn-arXiv, Citeseer and Wiki-CS, and three molecular graph classification datasets: BACE, BBBP and ClinTox. Statistics of these datasets are summarized in Table 7.
Cora McCallum et al. (2000) is a standard citation network dataset consisting of 2,708 research papers in the field of machine learning. These papers are categorized into seven classes. The title and abstract of each paper are utilized as the textual attributes of the nodes. The dataset includes 5,429 citation links, constructing a graph structure among the papers. The Cora dataset is commonly used for node classification and evaluating the performance of graph neural network models.
Citeseer Giles et al. (1998) is a citation network dataset comprising 3,312 research papers primarily from the fields of computer science and information technology. These papers are categorized into six classes based on their research areas. The title and abstract of each paper are used as the textual attributes of the nodes, providing a semantic representation of the content. The dataset includes 4,732 citation links, forming a graph structure that represents the citation relationships between the papers. The Citeseer dataset is widely used for node classification tasks, and it serves as a benchmark for evaluating the performance of graph neural network models.
PubMed Sen et al. (2008) is another citation network dataset containing 19,717 research papers from the biomedical field. These papers are categorized into three classes. The title and abstract of each paper are utilized as the textual attributes of the nodes. The dataset includes 44,338 citation links, constructing a graph structure among the papers. The PubMed dataset is also used for node classification tasks, particularly in testing large-scale graph neural network models.
ogbn-arXiv Hu et al. (2020) is part of the Open Graph Benchmark (OGB) and contains 169,343 academic papers scraped from arXiv. These papers are time-ordered by their submission dates and categorized into 40 subject areas. In this dataset, the nodes represent arXiv papers, and the edges represent citation relationships. The dataset includes 1,166,243 citation links, constructing a citation network among the papers. The ogbn-arXiv dataset is commonly used for node classification and studying time-sensitive graph neural network models.
Wiki-CS Mernyei and Cangea (2022) is derived from Wikipedia and consists of 11,701 web pages related to computer science topics. These pages are categorized into 10 classes, each representing a different area of computer science such as artificial intelligence, computer architecture, and software engineering. The text content of each page is used as the textual attributes of the nodes, capturing the thematic essence of the pages. The dataset includes 216,123 hyperlinks, which construct a graph structure connecting the web pages. The Wiki-CS dataset is often utilized for node classification tasks and for evaluating graph-based machine learning models.
BACE Wu et al. (2018) is a collection of inhibitors of human beta-secretase 1 (BACE-1) and provides both quantitative IC50 values and qualitative binary labels indicating the binding results. The BACE dataset comprises 1,513 molecular graphs for the molecular property prediction.
BBBP Wu et al. (2018) is used for predicting blood-brain barrier permeability (BBBP), which is crucial for determining whether a molecule can cross the blood-brain barrier. The BBBP dataset contains 2,039 molecular graphs for the binary graph classification task.
ClinTox Wu et al. (2018) is a collection of drugs approved by the FDA and those that have failed clinical trials due to toxicity reasons. The dataset encompasses two classification tasks for 1,491 drug molecules: (1) clinical trial toxicity and (2) FDA approval status.
For Cora, PubMed, Citeseer, BACE, BBBP and ClinTox, we randomly split nodes into 60%, 20%, and 20% for training, validation and testing, and measure the performance of all models on the test set.
For ogbn-arXiv we split the dataset as suggested in Hu et al. (2020).
For Wiki-CS, we split the dataset as suggested in Mernyei and Cangea (2022).
Table 7: Datasets statistics.
Dataset #Graphs Avg.# Nodes Avg.# Edges # Classes
Cora 1 2,708 5,429 7
PubMed 1 19,717 44,338 3
ogbn-arXiv 1 169,343 1,166,243 40
Citeseer 1 3,312 4,732 6
Wiki-CS 1 11,701 216,123 10
BACE 1,513 34.1 73.7 1
BBBP 2,039 23.9 51.6 1
ClinTox 1,491 26.1 55.5 2
Appendix B Baselines